# Fractionalized Prethermalization in a Driven Quantum Spin Liquid

### v2

@@ Abstract
Figure 3: Preparation of a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$. (a) Representation of a toric code state of effective spins $\bf{\tau}$ on the super lattice (dashed square) formed by the $z$-bonds of the Kitaev honeycomb model. The blue (yellow) diamonds indicate the vertical and horizontal sublattices on which the effective spins $\bf{\tau}$ live. (b) After applying a unitary transformation $U_{vh}$ on this toric code state, we obtain a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$. (c, d) The system in Eq. (2) initialized with $|\tilde{{\Psi}}_{0}\rangle$ exhibits similar nonequilibrium fractionalized dynamics as obtained with the ground state of Kitaev honeycomb model $|{\Psi}_{0}\rangle$. The simulations are performed for $V=1.0$ and $J=0.02$.
Experimental feasibility.—One intriguing prospect is to experimentally observe fractionalized prethermalization. First, we discuss the implementation of the dynamical Hamiltonian, which consists of three terms, the Kitaev honeycomb model $H_{\rm K}$, the Heisenberg interaction $H_{\rm J}$, and the three spin interaction $H_{\rm V}$. The first two can be directly decomposed into two-qubit Ising gates which in principle can be realized in various quantum architectures such as superconducting quantum processor [63, 64] and trapped atoms or molecules [65, 66, 67]. Moreover, the three spin term can also be conveniently prepared with two-qubit Ising gates by noting that $\sigma^{x}_{j}\sigma^{y}_{k}\sigma^{z}_{l}\propto(\sigma^{x}_{j}\sigma^{x}_{i})(\sigma^{y}_{k}\sigma^{y}_{i})(\sigma^{z}_{l}\sigma^{z}_{i})$. Second, the experimental preparation of $|\Psi_{0}\rangle$, the ground state of Kitaev honeycomb model, as an initial state is highly nontrivial. However, we need not to start in the ground state of the model, but a flux eigenstate at low energy density is sufficient. Thus, we propose a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$ which can be more easily realized in experiments and can lead to similar results as those obtained with $|\Psi_{0}\rangle$. Our proposal for preparing $|\tilde{{\Psi}}_{0}\rangle$ is motivated by the idea that the ground state of the Kitaev honeycomb model in the gapped $A$ phase is continuously connected to a toric code state [56]. We can, therefore, leverage previous work for the preparation o $|\tilde{{\Psi}}_{0}\rangle$, which showed that a toric code state can be efficiently prepared with a finite-depth quantum circuit [68].
@@ Acknowledgements.
[…]
${}^{1}$epartment of Physics, Technical University of Munich, 85748 Garching, Germany
${}^{2}$unich Center for Quantum Science and Technology (MCQST), Schellingstr. 4, 80799 Munich, Germany
${}^{3}$lackett Laboratory, Imperial College London, London SW7 2AZ, United Kingdom
@@ I Thermal expectation value of energy and flux
$H_{\rm K}(\{u_{ij}\})=\frac{i}{2}\left((c^{0}_{A})^{T},(c^{0}_{B})^{T}\right)\left(\begin{array}[]{cc}0&M\\ -M^{T}&0\end{array}\right)\left(\begin{array}[]{c}c^{0}_{A}\\ c^{0}_{B}\end{array}\right),$ (6)
@@
Hamiltonian in Eq. (6) can be diagonalized by singular-value decomposition (SVD) of the $N\times{}N$ matrix $M$ ($N$ the number of unit cell of a honeycomb lattice), $M=U\Lambda{}V^{T}$. Here $U^{T}U=I$, $V^{T}V=I$, and $\Lambda{}_{mn}=\delta_{nm}\epsilon_{m}$ is a semi-positive diagonal matrix. Then we can define new Majorana fermions as
$\begin{split}&(\alpha_{1},...,\alpha_{N})=(c^{0}_{A,1},...,c^{0}_{A,N})U,\\ &(\beta_{1},...,\beta_{N})=(c^{0}_{B,1},...,c^{0}_{B,N})V.\\ \end{split}$ (7)
@@
$H_{\rm K}(\{u_{ij}\})=i\sum_{m}\epsilon_{m}\alpha_{m}\beta_{m}=\sum_{m}\epsilon_{m}\left(2a^{\dagger}_{m}a_{m}-1\right),$ (8)
@@
By iterating over all possible flux configurations, the whole spectrum of $H_{\rm K}$ in the enlarged Majorana Hilbert space can be obtained according to Eq. (8). However, half of the eigenstates in the enlarged Hilbert are unphysical. When calculating the thermal expectation values, we should only account for the contribution from physical states.
Figure 4: Preparation of a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$. (a) Initialize the system with a product state $|\uparrow\rangle^{\otimes{}24}=|\Uparrow\rangle^{\otimes{}12}$. (b) Apply Hadamard gates $H$ on all representative effective spins. (c)-(e) Apply CNOT gates to obtain a toric code state in terms of effective spins. (f) Apply a unitary transformation $U_{vh}=\prod_{\rm vertical\ sites}U_{v}\prod_{\rm horizontal\ sites}U_{h}$. The unitaries $U_{v}$ are applied on the vertical sites (blue diamonds) and $U_{h}$ on the horizontal sites (yellow diamonds), see Eq. (19b). (g) This algorithm realizes a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$ for Kitaev honeycomb model on a surface.
@@
$\mathcal{P}=\prod_{j}\left(\frac{1+D_{j}}{2}\right).$ (9)
The effect of projector $\mathcal{P}$ is to annihilate unphysical states. In accordance with Ref. [75], $\mathcal{P}$ can be alternatively written as
$\mathcal{P}=\mathcal{S}\left(\frac{1+\mathcal{D}}{2}\right),$ (10)
@@
$Q=\left(\begin{array}[]{cc}0&U\\ V&0\end{array}\right),$ (13)
@@
$\mathcal{D}=(-1)^{\theta}{\rm det}(Q)(-1)^{N_{a}}\prod_{\langle{}ij\rangle_{a}}u_{ij},~{}$ (14)
where $N_{a}=\sum_{m}a^{\dagger}_{m}a_{m}$. In our case, e.g., a torus without any twisted boundaries, $\theta=L_{x}+L_{y}$ with $L_{x}$ and $L_{y}$ the length of a torus along $x$ and $y$ directions, respectively. By employing Eq. (14), all of the physical eigenstates can be selected.
@@
$Z(\beta)=\sum_{\{w_{p}\}}\sum_{E_{n}}\langle{}\psi_{n}(\{u_{ij}\})|\mathcal{P}e^{-\beta{}E_{n}}\mathcal{P}|\psi_{n}(\{u_{ij}\})\rangle$ (15)
@@
$\begin{split}&E(\beta)=\sum_{\{w_{p}\}}\sum_{E_{n}}\langle{}\psi_{n}(\{u_{ij}\})|\mathcal{P}E_{n}e^{-\beta{}E_{n}}\mathcal{P}|\psi_{n}(\{u_{ij}\})\rangle/Z(\beta),\\ &W(\beta)=\sum_{\{w_{p}\}}\sum_{E_{n}}\left\langle{}\psi_{n}(\{u_{ij}\})\left|\mathcal{P}{}\sum_{p}w_{p}e^{-\beta{}E_{n}}\mathcal{P}\right|\psi_{n}(\{u_{ij}\})\right\rangle/Z(\beta).\end{split}$ (16)
@@
$H=\left(\begin{array}[]{cccc}1/{\sqrt{2}}&&&1/{\sqrt{2}}\\ &1&&\\ &&1&\\ 1/{\sqrt{2}}&&&-1/{\sqrt{2}}\\ \end{array}\right),$ (17)
@@
$\begin{split}\mbox{CNOT}=&~{}|\Uparrow_{c}\Uparrow_{t}\rangle\langle\Uparrow_{c}\Uparrow_{t}|+|\Uparrow_{c}\Downarrow_{t}\rangle\langle\Uparrow_{c}\Downarrow_{t}|+\\ &~{}|\Downarrow_{c}\Uparrow_{t}\rangle\langle\Downarrow_{c}\Downarrow_{t}|+|\Downarrow_{c}\Downarrow_{t}\rangle\langle\Downarrow_{c}\Uparrow_{t}|+U_{\rm arb},\end{split}~{}$ (18)
where the subindex $c$ ($t$) stands for the control (target) two-site qubit and again $U_{\rm arb}$ is an arbitrary $12\times{}12$ unitary matrix spanned by $(|\uparrow_{c}\downarrow_{c}\rangle,|\downarrow_{c}\uparrow_{c}\rangle)\otimes{}(|\Uparrow_{t}\rangle,|\uparrow_{t}\downarrow_{t}\rangle,|\downarrow_{t}\uparrow_{t}\rangle,|\Downarrow_{t}\rangle)\oplus{}(|\Uparrow_{c}\rangle,|\Downarrow_{c}\rangle)\otimes(|\uparrow_{t}\downarrow_{t}\rangle,|\downarrow_{t}\uparrow_{t}\rangle)$. Note that the CNOT gate in Eq. (18) involves four physical qubits.
Figure 5: Mechanism for the heating of the flux degree of freedom. (a) The dynamics of renormalized flux on short time scales for various drive frequencies $\omega$. (b) The same observable but with fixed $\omega=4$. The inset demonstrates a collapse of the short time flux dynamics $W(t)$ when times are rescaled by $J^{2}$. The simulations are performed for $V=1.0$ on a $3\times{}4$ torus.
@@
$U_{v}=\left(\begin{array}[]{cccc}(1+i)/2&&&(1+i)/2\\ &1&&\\ &&1&\\ (-1+i)/2&&&(1-i)/2\\ \end{array}\right),$ (19a)
@@
$U_{h}=\left(\begin{array}[]{cccc}e^{-i\pi/4}&&&0\\ &1&&\\ &&1&\\ &&&e^{i\pi/4}\\ \end{array}\right).$ (19b)
Similar to the Hadmard gates, the $2\times 2$ identity submatrices in Eq. (19b) can be arbitrary $2\times 2$ unitary matrices.
@@ III Mechanism of short-time flux heating
In order to examine which mechanism heats the flux sector, we vary the Heisenberg coupling $J$ and study the dynamics of $W(t)$. When rescaling the time with $J^{2}$ as $t\rightarrow{}tJ^{2}$, the flux dynamics $W(t)$ give a collapse of the short time data for various $J$, see Fig. 5. This data collapse can be explained from the following arguments. Denoting the initial state (a zero-flux state) as $|\psi_{0}\rangle$, the short-time dynamics of plaquette operator $W_{p}$ up to the second order of $J$ reads
$W(t)\sim\langle\psi_{0}|(1+iTJ\tilde{H}_{\rm J})W_{p}(1-iTJ\tilde{H}_{\rm J})|\psi_{0}\rangle,$ (20)
@@
$\frac{{\rm d}W(t)}{{\rm d}t}\sim{}J^{2}\langle\psi_{0}|H_{\rm J}W_{p}H_{\rm J}|\psi_{0}\rangle.$ (21)

### v4

@@ Abstract
Figure 3: Preparation of a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$. (a) Representation of a toric code state of effective spins $\bf{\tau}$ on the super lattice (dashed square) formed by the $z$-bonds of the Kitaev honeycomb model. The blue (yellow) diamonds indicate the horizontal (vertical) sublattices on which the effective spins $\bf{\tau}$ live. (b) After applying a unitary transformation $U_{ab}$ on the toric code state, we obtain a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$. (c, d) The system in Eq. (2) initialized with $|\tilde{{\Psi}}_{0}\rangle$ exhibits similar nonequilibrium fractionalized dynamics as obtained with the ground state of Kitaev honeycomb model $|{\Psi}_{0}\rangle$. The simulations are performed for $V=1.0$ and $J=0.02$.
Experimental feasibility.—One intriguing prospect is to experimentally observe fractionalized prethermalization. First, we discuss the implementation of the dynamical Hamiltonian, which consists of three terms, the Kitaev honeycomb model $H_{\rm K}$, the Heisenberg interaction $H_{\rm J}$, and the three spin interaction $H_{\rm V}$. The first two can be directly decomposed into two-qubit Ising gates which in principle can be realized in various quantum architectures such as superconducting quantum processor [63, 64] and trapped atoms or molecules [65, 66, 67]. Moreover, the three spin term can also be conveniently prepared with two-qubit Ising gates by noting that $\sigma^{x}_{j}\sigma^{y}_{k}\sigma^{z}_{l}\propto(\sigma^{x}_{j}\sigma^{x}_{i})(\sigma^{y}_{k}\sigma^{y}_{i})(\sigma^{z}_{l}\sigma^{z}_{i})$. Second, the experimental preparation of $|\Psi_{0}\rangle$, the ground state of Kitaev honeycomb model, as an initial state is highly nontrivial. However, we need not to start in the ground state of the model, but a flux eigenstate at low energy density is sufficient. Thus, we propose a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$ which can be more easily realized in experiments and can lead to similar results as those obtained with $|\Psi_{0}\rangle$. Our proposal for preparing $|\tilde{{\Psi}}_{0}\rangle$ is motivated by the idea that the ground state of the Kitaev honeycomb model in the gapped $A$ phase is continuously connected to a toric code state [56]. We can, therefore, leverage previous work for the preparation of $|\tilde{{\Psi}}_{0}\rangle$, which showed that a toric code state can be efficiently prepared with a finite-depth quantum circuit [68].
@@ Acknowledgements.
Hui-Ke Jin1, Johannes Knolle1,2,3, Michael Knap1,2
1Department of Physics, Technical University of Munich, 85748 Garching, Germany
2Munich Center for Quantum Science and Technology (MCQST), Schellingstr. 4, 80799 Munich, Germany
3Blackett Laboratory, Imperial College London, London SW7 2AZ, United Kingdom
@@ I Thermal expectation value of energy and flux
$H_{\rm K}(\{u_{ij}\})=\frac{i}{2}\left((c^{0}_{A})^{T},(c^{0}_{B})^{T}\right)\left(\begin{array}[]{cc}0&M\\ -M^{T}&0\end{array}\right)\left(\begin{array}[]{c}c^{0}_{A}\\ c^{0}_{B}\end{array}\right),$ (S3)
@@
Hamiltonian in Eq. (S3) can be diagonalized by singular-value decomposition (SVD) of the $N\times{}N$ matrix $M$ ($N$ the number of unit cell of a honeycomb lattice), $M=U\Lambda{}V^{T}$. Here $U^{T}U=I$, $V^{T}V=I$, and $\Lambda{}_{mn}=\delta_{nm}\epsilon_{m}$ is a semi-positive diagonal matrix. Then we can define new Majorana fermions as
$\begin{split}&(\alpha_{1},...,\alpha_{N})=(c^{0}_{A,1},...,c^{0}_{A,N})U,\\ &(\beta_{1},...,\beta_{N})=(c^{0}_{B,1},...,c^{0}_{B,N})V.\\ \end{split}$ (S4)
@@
$H_{\rm K}(\{u_{ij}\})=i\sum_{m}\epsilon_{m}\alpha_{m}\beta_{m}=\sum_{m}\epsilon_{m}\left(2a^{\dagger}_{m}a_{m}-1\right),$ (S5)
@@
By iterating over all possible flux configurations, the whole spectrum of $H_{\rm K}$ in the enlarged Majorana Hilbert space can be obtained according to Eq. (S5). However, half of the eigenstates in the enlarged Hilbert are unphysical. When calculating the thermal expectation values, we should only account for the contribution from physical states.
Figure S1: Preparation of a zero-flux state $|\tilde{{\Psi}}_{0}\rangle$. (a) Initialize the system with a product state $|\uparrow\rangle^{\otimes{}24}=|\Uparrow\rangle^{\otimes{}12}$. (b) Apply Hadamard gates $H$ on all representative effective spins. (c)-(e) Apply CNOT gates to obtain a toric code state in terms of effective spins. (f) Apply a unitary transformation ${U}_{ab}$, which consists of unitaries $U_{a}$ and $U_{b}$ applied on the horizontal and vertical sites, respectively. (g) This algorithm realizes a zero-flux state for the Kitaev honeycomb model on a surface.
@@
$\mathcal{P}=\prod_{j}\left(\frac{1+D_{j}}{2}\right).$ (S6)
The effect of $\mathcal{P}$ is to annihilate unphysical states. In accordance with Ref. [75], $\mathcal{P}$ can be alternatively written as
$\mathcal{P}=\mathcal{S}\left(\frac{1+\mathcal{D}}{2}\right),$ (S7)
@@
$Q=\left(\begin{array}[]{cc}0&U\\ V&0\end{array}\right),$ (S10)
@@
$\mathcal{D}=(-1)^{\theta}{\rm det}(Q)(-1)^{N_{a}}\prod_{\langle{}ij\rangle_{a}}u_{ij},~{}$ (S11)
where $N_{a}=\sum_{m}a^{\dagger}_{m}a_{m}$. In our case, e.g., a torus without any twisted boundaries, $\theta=L_{x}+L_{y}$ with $L_{x}$ and $L_{y}$ the length of a torus along $x$ and $y$ directions, respectively. By employing Eq. (S11), all of the physical eigenstates can be selected.
@@
$Z(\beta)=\sum_{\{w_{p}\}}\sum_{E_{n}}\langle{}\psi_{n}(\{u_{ij}\})|\mathcal{P}e^{-\beta{}E_{n}}\mathcal{P}|\psi_{n}(\{u_{ij}\})\rangle$ (S12)
@@
$\begin{split}&E(\beta)=\sum_{\{w_{p}\}}\sum_{E_{n}}\langle{}\psi_{n}(\{u_{ij}\})|\mathcal{P}E_{n}e^{-\beta{}E_{n}}\mathcal{P}|\psi_{n}(\{u_{ij}\})\rangle/Z(\beta),\\ &W(\beta)=\sum_{\{w_{p}\}}\sum_{E_{n}}\left\langle{}\psi_{n}(\{u_{ij}\})\left|\mathcal{P}{}\sum_{p}w_{p}e^{-\beta{}E_{n}}\mathcal{P}\right|\psi_{n}(\{u_{ij}\})\right\rangle/Z(\beta).\end{split}$ (S13)
@@
$H=\left(\begin{array}[]{cccc}1/{\sqrt{2}}&&&1/{\sqrt{2}}\\ &1&&\\ &&1&\\ 1/{\sqrt{2}}&&&-1/{\sqrt{2}}\\ \end{array}\right),$ (S14)
@@
$\begin{split}\mbox{CNOT}=&~{}|\Uparrow_{c}\Uparrow_{t}\rangle\langle\Uparrow_{c}\Uparrow_{t}|+|\Uparrow_{c}\Downarrow_{t}\rangle\langle\Uparrow_{c}\Downarrow_{t}|+\\ &~{}|\Downarrow_{c}\Uparrow_{t}\rangle\langle\Downarrow_{c}\Downarrow_{t}|+|\Downarrow_{c}\Downarrow_{t}\rangle\langle\Downarrow_{c}\Uparrow_{t}|+U_{\rm arb},\end{split}~{}$ (S15)
where the subindex $c$ ($t$) stands for the control (target) two-site qubit and again $U_{\rm arb}$ is an arbitrary $12\times{}12$ unitary matrix spanned by $(|\uparrow_{c}\downarrow_{c}\rangle,|\downarrow_{c}\uparrow_{c}\rangle)\otimes{}(|\Uparrow_{t}\rangle,|\uparrow_{t}\downarrow_{t}\rangle,|\downarrow_{t}\uparrow_{t}\rangle,|\Downarrow_{t}\rangle)\oplus{}(|\Uparrow_{c}\rangle,|\Downarrow_{c}\rangle)\otimes(|\uparrow_{t}\downarrow_{t}\rangle,|\downarrow_{t}\uparrow_{t}\rangle)$. Note that the CNOT gate in Eq. (S15) involves four physical qubits.
Figure S2: Mechanism for the heating of the flux degree of freedom. (a) The dynamics of renormalized flux on short time scales for various drive frequencies $\omega$. (b) The same observable but with fixed $\omega=4$. The inset demonstrates a collapse of the short time flux dynamics $W(t)$ when times are rescaled by $J^{2}$. The simulations are performed for $V=1.0$ on a $3\times{}4$ torus.
@@
$U_{a}=\left(\begin{array}[]{cccc}(1+i)/2&&&(1+i)/2\\ &1&&\\ &&1&\\ (-1+i)/2&&&(1-i)/2\\ \end{array}\right),$ (S16a)
@@
$U_{b}=\left(\begin{array}[]{cccc}e^{-i\pi/4}&&&0\\ &1&&\\ &&1&\\ &&&e^{+i\pi/4}\\ \end{array}\right).$ (S16b)
Similar to the Hadmard gates, the $2\times 2$ identity submatrices in Eq. (S16) can be arbitrary $2\times 2$ unitary matrices.
@@ III Mechanism of short-time flux heating
In order to examine which mechanism heats the flux sector, we vary the Heisenberg coupling $J$ and study the dynamics of $W(t)$. When rescaling the time with $J^{2}$ as $t\rightarrow{}tJ^{2}$, the flux dynamics $W(t)$ give a collapse of the short time data for various $J$, see Fig. S2. This data collapse can be explained from the following arguments. Denoting the initial state (a zero-flux state) as $|\psi_{0}\rangle$, the short-time dynamics of plaquette operator $W_{p}$ up to the second order of $J$ reads
$W(t)\sim\langle\psi_{0}|(1+iTJ\tilde{H}_{\rm J})W_{p}(1-iTJ\tilde{H}_{\rm J})|\psi_{0}\rangle,$ (S17)
@@
$\frac{{\rm d}W(t)}{{\rm d}t}\sim{}J^{2}\langle\psi_{0}|H_{\rm J}W_{p}H_{\rm J}|\psi_{0}\rangle.$ (S18)